Two resistors have resistances R1 and R2. When the resistors are connected in series to a 11.1-V battery, the current from the battery is 1.96 A. When the resistors are connected in parallel to the battery, the total current from the battery is 9.16 A. DetermineR1 and R2. (Enter your answers from smallest to largest.)

R1 = ?
R2 = ?
Answer #1

Let us assume that internal resistance of battery is zero.
11.1/(R1 + R2) = 1.96 or R1+R2 = 5.66 ohm
11.1/(R1R2/R1+R2) = 9.16 or 11.1(R1+R2)/R1R2 = 9.16 substituting for R1 + R2 we have
R1R2 = 11.1*5.66/9.16 = 6.796 or 4R1R2 = 27.19
(R2 - R1)^2 = (R2+R1)^2 - 4R1R2 = 32.035 - 27.19 = 4.845 or
R2 - R1 = 2.20 or 2R2 = 7.86 or R2 =3.93 ohm and R1 = 1.73 ohm